OpenStudy (anonymous):

prove the identity: tan(x)/(sin(x)-cos(x)) = (sin^2(x) + sin(x)cos(x))/(cos(x)-2cos^3(x))

OpenStudy (anonymous):

i. hate. trig. identities.

OpenStudy (anonymous):

that "2" in \[\cos^2(x)-2\cos^3(x)\] is really messing things up \[\frac{\frac{b}{a}}{b-a}=\frac{b^2+ab}{a-2a^3}\] \[\frac{b}{a(b-a)}=\frac{b(b+a)}{a(1-2a^2)}\] i can't figure out what to do with the 2!

OpenStudy (anonymous):

multiply both sides by (sinx-cosx) to get tanx = sinx/cosx = (sin^3x -cos^2xsinx)/(cosx-2cos^3x) =sinx(Sin^2x-Cos^2x)/cosx(1-2cos^2x) and the bracketed terms are both equal to Cos2x QED

OpenStudy (anonymous):

Sorry, forgot the sign, both equal to - Cos2x

OpenStudy (anonymous):

You happy with that?

OpenStudy (anonymous):

I don't completely follow.

OpenStudy (anonymous):

Which part?

OpenStudy (anonymous):

Just a sec.

OpenStudy (anonymous):

I shortened it a bit at the beginning because I didn't want have to type it all out.

OpenStudy (anonymous):

I'm attaching a pdf with my attempt to use your solution. Can you point out where I've gone wrong?

OpenStudy (anonymous):

I think I might know where I messed up. I'm trying again.

OpenStudy (anonymous):

The first step, "multiply both sides by (sinx-cosx) to get" this is just to get tan x on the left hand side by itself. Then multiply out the top row. Then factor sin out of the top and cos out of the bottom.

OpenStudy (anonymous):

very nice.

OpenStudy (anonymous):

I don't think the first step of your proof is legal. I'm supposed to take one side and transform it into the other side.

OpenStudy (anonymous):

You're first step assumes that they're equal from the start. Then there's no reason to prove the identity.

OpenStudy (anonymous):

My first step hasn't changed anything at all, it's the same thing you wrote rearranged. Besides that, it is perfectly OK in a proof to make an assumption and then see if you get a contradiction. If the two sides had not been equal, then I would not have been able to show equality as I did.

OpenStudy (anonymous):

I think maybe you are assuming that that is what you have to do. Probably there is a way, you can try now that you have seen the possibilities. If all you were asked to do is prove equality, then my method is 100% satisfactory.

OpenStudy (anonymous):

All I know is that if a properly trained teacher marked it wrong, he should be replaced.

OpenStudy (anonymous):

How does 1-2cos^2x = -cos 2x? Haven't learned that identity yet.

OpenStudy (anonymous):

Are you sure? Cos 2x = cos^x - sin^x (sum formula). Then use cos^2x + sin^2x =1 (Pythagorean identity)

OpenStudy (anonymous):

Thanks. Sum and difference formulas are in the next chapter. There should be a way to do this proof without the -cos 2x cancellation. I'm still working on it.

OpenStudy (anonymous):

OK, I did it your way (what a pain). Start off the same way, get rid of sin/cos (for both of them) and you are left with 1/(sinx-cosx) = (sinx + cosx)/(1-2cos^2x) (I'm just using = for convenience) Now multiply the top and bottom of the lhs by (sinx + cosx) to get rid of that (for both) leaving to show 1/((sin x-cosx)(sinx+cosx)) = 1/(1-2cos^2x) and I will leave it to you to show that both denominators are the same. (no double angles, basic operations only, no intermingling, hope I haven't broken any more of your teacher's rules!)

OpenStudy (anonymous):

Thanks for the help. I discovered a solution last night by factoring by grouping, working with the right hand side. It's attached. I'm still using your trick of multiplying top and bottom by sin x - cos x.